Write the first five terms of the sequence in which the nth term is 5n-4.
Answer by amarjeeth123(575)

The nth term is 5n-4
The first term is 5(1)-4=1
The second term is 5(2)-4=6
The third term is 5(3)-4=11
The fourth term is 5(4)-4=16
The fifth term is 5(5)-4=21
The first five terms of the series are 1,6,11,16,21.